OpenStudy (anonymous):

A circle has a diameter with endpoints (-7, -1) and (-5, -9). What is the equation of the circle? r2 = (x + 6)2 + (y + 5)2 r2 = (x + 4)2 + (y + 1)2 r2 = (x + 6)2 + (y - 1)2 r2 = (x + 4)2 + (y + 5)2

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

center of the circle is a midpoint of the two given end points. so center of the circle is \[\left(\frac{-7+-5}{2},\frac{-1+-9}{2}\right) = (-6, -5)\]

OpenStudy (anonymous):

that makes sense

OpenStudy (anonymous):

Hence equation of the circle is \[(x+6)^2+(y+5)^2=r^2\]

OpenStudy (anonymous):

Why does it change from negative to positive?

OpenStudy (anonymous):

because the formula is (x - h)^2 + (y - k)^2 = r^2, so when you put negative numbers for h and k, it becomes positive

OpenStudy (anonymous):

Ohhh! That makes so much sense! I actually just went over that a few days ago! Can't believe I forgot. Wow! Thanks so much! :)

OpenStudy (anonymous):

Glad to help you :D